An 80N force causes a spring to compress 0.15m. What is the spring constant? What is the potential energy of the spring?


1
Expert's answer
2021-05-05T19:05:55-0400

(a) We can find the spring constant from the Hooke's Law:


"F=kx,""k=\dfrac{F}{x}=\dfrac{80\ N}{0.15\ m}=533\ \dfrac{N}{m}."

(b) We can find the potential energy of the spring as follows:


"PE=\dfrac{1}{2}kx^2,""PE=\dfrac{1}{2}\cdot533\ \dfrac{N}{m}\cdot(0.15\ m)^2=6.0\ J."